OpenStudy (anonymous):

A scientist needs 32 gallons of a 12% saline solution to complete an experiment. All he has in the lab are two liquids. One has only 7% saline while the other has 23% saline. How much of each should she mix together in order to get 32 gallons of12% solution.

OpenStudy (anonymous):

\[23x+7(32-x)=32\times 12\] should do it., if \(x\) represents the number of gallons of 23% solution

hero (hero):

x = gallons of 7% solution y = gallons of 23% solution x + y = 32 .07x + .23y = .12*32

OpenStudy (anonymous):

Thank you...

hero (hero):

Thank you... to whom?

OpenStudy (anonymous):

To both of yous... :-)

hero (hero):

You can only use one set of formulas bro.

OpenStudy (anonymous):

Okay thanks. I needed an idea how to start. Im getting a tutor because the school I went to before didnt prepare me for this, but i really like this site. Its all new I just have to learn how to use it better.